OpenStudy (anonymous):

F(x)=^3sqrt(x-1) +4 Find f^-1(x) ........ HELP!!!!

OpenStudy (asnaseer):

is this the correct equation?\[f(x)=3\sqrt{x-1}+4\]

OpenStudy (anonymous):

asnaseer if u could help me with my latest probs that would be awesome man!

OpenStudy (anonymous):

yea

OpenStudy (asnaseer):

ok, then what you need to do is invert this so you end up with x = ...

OpenStudy (asnaseer):

you can do this steps

OpenStudy (asnaseer):

"in steps"

OpenStudy (anonymous):

so u end up with [x=\sqrt[3]{x-1}+4\]

OpenStudy (anonymous):

for x and the same for y

OpenStudy (asnaseer):

no - let me try and do it step-by-step for you

OpenStudy (anonymous):

alright thank you

OpenStudy (asnaseer):

\[f(x)=3\sqrt{x-1}+4\]take 4 away from both sides to get:\[f(x)-4=3\sqrt{x-1}\]square both sides to get:\[(f(x)-4)^2=(3\sqrt{x-1})^2=9(x-1)\]divide both sides by 9 to get:\[\frac{(f(x)-4)^2}{9}=x-1\]add 1 to both sides to get:\[\frac{(f(x)-4)^2}{9}+1=x\]finally swap the left-hand and right-hand sides to get the answer:\[x=\frac{(f(x)-4)^2}{9}+1\]the usual practice is to then replace x on the LHS with f-1(x) and replace the f(x) on the RHS with x to get the inverse function:\[f^{-1}(x)=\frac{(x-4)^2}{9}+1\]

OpenStudy (asnaseer):

hope that makes sense.

OpenStudy (anonymous):

thank you!! :))

OpenStudy (asnaseer):

no problem - glad to be of service :-)